# Leveraging Quantum Superposition to Infer the Dynamic Behavior of a Spatial-Temporal Neural Network Signaling Model

### v2

@@ 4.3.2 Diffusion operator and search space
Here, $I$ is the $N\times N$ identity matrix, and $N=|\mathbb{S}_{n}|=2^{n}$. The diffusion operator amplifies the amplitudes of marked states identified by $U_{f}$, inverting their amplitudes about the average. In our context, Grover’s algorithm iteratively applies $U_{f}$ and $D$ to amplify the amplitudes of $s_{i}\in\mathbb{S}_{n}$ that are present in $\mathbb{V}$, which would allow efficient identification of elements in $S_{n}$.
@@ 4.3.3 Iterative application of Grover’s algorithm
where $\mu$ is the number of occurrences of $s_{i}$ in $\mathbb{V}$. The algorithm’s efficiency improves with higher $\mu$ values because fewer iterations are required. For real physically constructable networks, there are likely other physical and functional constraints. So optimizing $k$ may benefit from structural insights into $G(V,E)$, which could suggest and constrain patterns in $\mathbb{V}$. For example, signaling conditions that constrain the possible range of values $\Sigma_{r}$ can take.
@@
4.3.4 Probabilistic nature and verification
If the algorithm identifies $s_{i}\in\mathbb{V}$ once measured, we assume it is part of $S_{n}$. If the algorithm returns $s_{j}\neq s_{i}$, we conditionally assume $s_{i}\notin\mathbb{V}$. However, there is a non-zero chance of error because of the probabilistic nature of quantum mechanics. In a practical sense, increasing the confidence of set membership may require re-running the algorithm for the same $s_{i}$.
Once $S_{n}$ is identified, determining $S_{n}^{c}$ is straightforward. It is simply the complement of $S_{n}$ in $\mathbb{S}_{n}$: $S_{n}^{c}=\mathbb{S}_{n}\setminus S_{n}$. This allows us to construct both subsets efficiently using Grover’s algorithm, setting up the computations we will do in the Deutsch–Jozsa computations.
@@ 5 Evaluating Network Dynamics with the Deutsch-Jozsa Algorithm
With $S_{n}$ and $S_{n}^{c}$ defined, we can now use the Deutsch–Jozsa algorithm to determine if the network can sustain inherent recurrent activity beyond $T_{o}$, or if its dynamics will necessarily cease. However, as discussed above, the evaluation must be done on $S_{n}$, a subset of $\mathbb{S}_{n}=\{0,1\}^{n}$, rather than on the full domain.
To address this subset evaluation requirement, we extend $S_{n}$ into $\mathbb{S}_{n}$ by taking advantage of the relationship $S_{n}\cup S_{n}^{c}=\mathbb{S}_{n}$. This construction will let us interpret the Deutsch–Jozsa algorithm’s output to indirectly infer the behavior of $f(S_{n})$, as defined by equation 16. Here, $f(S_{n})$ refers to the input function into the Deutsch–Jozsa algorithm, which, as required, is evaluated across the full domain $\mathbb{S}_{n}$. Specifically, the elements of $S_{n}$ encode the key information about the network dynamics in the form of the node outputs at $T_{o}$ (equation 16). As such, in our implementation here, $f(S_{n})$ does not explicitly evaluate the elements of the input set $S_{n}$, but instead, it preserves them unchanged. It effectively acts as an identify function.
For elements in $S_{n}^{c}$, $f$ does not evaluate their values explicitly. Instead, all elements of $S_{n}^{c}$ are assigned a fixed value (either 0 or 1) for one algorithm run and then reassigned the other fixed value for a second subsequent run. This assignment ensures that we can infer the behavior of $f(S_{n})$ based on the algorithm’s output. To do this, we need to construct and evaluate a two-part unitary function $U_{f}$ across two distinct runs of the algorithm.
@@ 5.1.1 Behavior for $S_{n}$
For $s_{i}\in S_{n}$, the oracle function evaluates $f(s_{i})$, flipping the phase of |$\ket{s_{i}}$ when $f(s_{i})=1$ and leaving it unchanged when $f(s_{i})=0$, such that
$\text{If }f(s_{i})=1:\quad U_{f}\ket{s_{i}}\ket{-}=-\ket{s_{i}}\ket{-},$ (26a)
$|.\displaystyle\text{If }f(s_{i})=0:\quad U_{f}\ket{s_{i}}\ket{-}=\ket{s_{i}}\ket{-}.$ (26b)
@@ 5.1.2 Behavior for $S_{n}^{c}$
For $s_{i}\in S_{n}^{c}$, the oracle does not evaluate $f(s_{i})$ directly. Instead:
@@
    For the first run of the algorithm, all elements $s_{i}\in S_{n}^{c}$ are assigned $f(s_{i})=0$.
@@
    In a separate second run, all elements $s_{i}\in S_{n}^{c}$ are assigned $f(s_{i})=1$.
This ensures that $S_{n}^{c}$ does not contribute to the dynamic information encoded in $f(S_{n})$. But it satisfies the technical requirement that the input function $f(\cdot)$ operates over the entire domain $\mathbb{S}_{n}$ by filling in the values of $\mathbb{S}_{n}$ not encountered in $S_{n}$ that encodes the dynamics.
@@
$\text{If assigned }f(s_{i})=1:\quad U_{f}\ket{s_{i}}\ket{-}=-\ket{s_{i}}\ket{-},$ (27a)
$|.\displaystyle\text{If assigned }f(s_{i})=0:\quad U_{f}\ket{s_{i}}\ket{-}=\ket{s_{i}}\ket{-}.$ (27b)
@@ 5.2 Constructing $U_{f}$ to Evaluate $f(S_{n})$
    Evaluation of $S_{n}$: Apply $f(s_{i})$ to elements in $S_{n}$ according to equation 26.
@@
The second step evaluates $f(S_{n})$ by comparing each $\Sigma_{r}$ (encoded in $s_{i}$) with the threshold $\Sigma_{T}$. We implement this as a binary subtraction $\Sigma_{r}-\Sigma_{T}$, with the ancillary qubit encoding the result
@@
    |$\ket{1}$ if $\Sigma_{r}\geq\Sigma_{T}$, corresponding to $f(s_{i})=1$.
@@
    |$\ket{0}$ if $\Sigma_{r}<\Sigma_{T}$, corresponding to $f(s_{i})=0$.
@@ 5.2.1 Example of a Binary Subtraction for Evaluating $f(S_{n})$
Let’s work through a specific example of a binary subtraction operation. Consider a vertex $v_{j}\in G(V,E)$ with a corresponding entry $\rho_{i}\in\mathbb{V}$ that encodes the running summation value $\Sigma_{r}$ at the observation time $T_{o}$ for the specific vertex $v_{j}$. Assume a 7-bit representation of $\Sigma_{r}$ and the threshold $\Sigma_{T}$ to two-decimal precision, as discussed in Section 4.1.
@@
where |$\ket{\rho}$ encodes $\Sigma_{r}=0.85$ in binary, while |$\ket{\tau}$ encodes $\Sigma_{T}=0.50$. The ancillary qubit, initially in state |$\ket{0}$, will store the result of comparing $\Sigma_{r}\geq\Sigma_{T}$.
@@
In a physical quantum circuit this comparison can be implemented using a sequence of controlled gates that
@@
    Compare corresponding bits $\rho_{i}$ and $\tau_{i}$ sequentially from the most significant to least significant.
@@
    Stop and set the ancillary qubit when the first inequality $\rho_{i}>\tau_{i}$ is detected.
This binary subtraction comparison is efficient because it focuses only on detecting the sign of the result (i.e., positive or negative) rather than computing the full numerical value of $\Sigma_{r}-\Sigma_{T}$. Once $f(s_{i})$ is computed for all $s_{i}\in S_{n}$, the Deutsch–Jozsa algorithm evaluates whether $f(S_{n})$ is constant or not across the entire subset.
5.3 Measurement Interpretations of $U_{f}$
By running the Deutsch–Jozsa algorithm twice, assigning all elements of $S_{n}^{c}$ first to $f(s_{i})=0$ and then to $f(s_{i})=1$, measurement outcomes provide an interpretation that answers our question about the network dynamics:
  1. i.
    Quiescence: If $S_{n}^{c}=0$ and $U_{f}$ evaluates to a constant state |$\ket{0}^{\otimes n}$, the network is quiescent, because $f(S_{n})=0$ for all $s_{i}\in S_{n}$.
  2. ii.
    Epileptic State: If $S_{n}^{c}=1$ and $U_{f}$ evaluates to a constant state |$\ket{1}^{\otimes n}$, the network is epileptic, because $f(S_{n})=1$ for all $s_{i}\in S_{n}$.
@@ 5.3 Measurement Interpretations of $U_{f}$
    Dynamic Activity: If $U_{f}$ is not constant for either case, we interpret this result as the network being able to potentially maintain inherent dynamic activity due to the measurement outcome reflecting what must be a mix of active and inactive vertices in $S_{n}$.
@@ 6 Comparative Algorithmic Complexity
This last section analyzes the algorithmic complexity of solving the problem using the Grover and Deutsch–Jozsa algorithms versus a purely classical approach. We will show that using these quantum algorithms conveys a significant computational advantage and speedup and demonstrate how such an advantage scales, i.e., becomes greater, as the size of the network being analyzed increases.
@@ 6.1 Classical complexity
    Evaluate $F(V,T_{o})=f_{j}(T_{o})$ for each $v_{j}\in V$ in order to compare $\Sigma_{r}$ and $\Sigma_{T}$ (equation 16). This requires iterating over all vertices.
@@
    Combine the results of $f_{j}(T_{o})$ for all $v_{j}\in V$ to determine whether the network is quiescent, epileptic, or dynamically active.
@@ 6.2 Quantum complexity: Grover’s algorithm
Grover’s algorithm provides a quadratic speedup for searching an unstructured set. To determine $S_{n}$, we use Grover’s algorithm to find all $s_{i}\in\mathbb{S}_{n}$ that belong to $S_{n}$. Let $\mu$ denote the number of occurrences of a particular $s_{i}$ in $\mathbb{V}$, and $|\mathbb{S}_{n}|=2^{n}$. The number of iterations required for Grover’s algorithm to amplify the amplitude of each solution is approximately
@@ 6.3 Quantum complexity: Deutsch–Jozsa Algorithm
The Deutsch–Jozsa algorithm evaluates whether $f(S_{n})$ is constant or not across all elements of $\mathbb{S}_{n}$. Unlike a classical approach that would require checking each $s_{i}\in\mathbb{S}_{n}$, the Deutsch–Jozsa algorithm achieves this in a constant number of evaluations:
@@ 6.4 Comparative analysis
As such, the task of evaluating $f(S_{n})$ for all vertices in $V$ can be summarized by
@@
For large $N$ and $n$, a quantum algorithmic approach provides a significant increase in computational efficiency. Grover’s algorithm provides a quadratic speedup in $2^{n}$, while Deutsch–Jozsa evaluates $f(S_{n})$ in constant time.
@@
Then subsequently evaluating $f(S_{n})$ with Deutsch–Jozsa adds a constant overhead of $O_{\text{Deutsch-Jozsa}}=O(1).$
@@
We can assess the comparative speedup advantage of the quantum approach versus the classical approach by computing the ratios of their respective complexities. The classical approach needs $O(7\times 10^{6})$, while the quantum approach requires $O(1131)$. This then translates to a speedup factor of approximately
@@
By ‘faster’, we specifically mean that a quantum approach, as devised here, requires significantly fewer computational steps to arrive at the result. This reduction translates into less computational time, assuming similar hardware clock speeds and negligible algorithmic overhead and assuming a physical quantum computer with sufficiently stable logical qubits capable of encoding a circuit that encodes the construction we outline in this work.

### v4

@@ 4.3.2 Diffusion operator and search space
Here, $I$ is the $N\times N$ identity matrix, and $N=|\mathbb{S}_{n}|=2^{n}$. The diffusion operator amplifies the amplitudes of marked states identified by $U_{f}$, inverting their amplitudes about the average. In our context, Grover’s algorithm iteratively applies $U_{f}$ and $D$ to mark $s_{i}\in\mathbb{S}_{n}$ that are present in $\mathbb{V}_{n}$, which will allow efficient identification of elements in $S_{n}$ (see the discussion below in Section 7.2).
@@ 4.3.3 Iterative application of Grover’s algorithm
where $\mu$ is the number of occurrences of $s_{i}$ in $\mathbb{V}_{n}$. The algorithm’s efficiency improves with higher $\mu$ values because fewer iterations are required. For real physically constructable networks, there are likely other physical and functional constraints. So optimizing $k$ may benefit from structural insights into $G(V,E)$, which could suggest and constrain patterns in $\mathbb{V}_{n}$. For example, signaling conditions that constrain the possible range of values $\Sigma_{r}$ can take. In practicality, though, while we can successfully mark the states $s_{i}\in S_{n}$, amplification is more complex than this because we are applying Grover to multiple states simultaneously. Again, see the detailed discussion in in Section 7.2 below in the paper.
@@
Once $S_{n}$ is identified and marked, determining $S_{n}^{c}$ is straightforward. It is simply the complement of $S_{n}$ in $\mathbb{S}_{n}$: $S_{n}^{c}=\mathbb{S}_{n}\setminus S_{n}$. This allows us to construct both subsets efficiently using Grover, setting up the computations we will do in the Deutsch–Jozsa computations.
@@ 5 Evaluating Network Dynamics with the Deutsch-Jozsa Algorithm
With $S_{n}$ and $S_{n}^{c}$ defined, we can now use the Deutsch–Jozsa algorithm to determine if the network can sustain inherent recurrent activity beyond $T_{o}$, or if its dynamics will necessarily cease. However, as emphasized in Box 2., the evaluation must be done on $S_{n}$, a subset of $\mathbb{S}_{n}=\{0,1\}^{n}$, rather than on the full domain.
To address this subset evaluation requirement, we extend $S_{n}$ into $\mathbb{S}_{n}$ by taking advantage of the relationship $S_{n}\cup S_{n}^{c}=\mathbb{S}_{n}$. This construction will let us interpret the Deutsch–Jozsa algorithm’s output to indirectly infer the behavior of $\mathcal{T}(S_{n})$. Specifically, the elements of $S_{n}$ encode the key information about the network dynamics.
For elements in $S_{n}^{c}$, there is no explicit evaluation. Instead, all elements of $S_{n}^{c}$ are assigned a fixed value (either 0 or 1) for one algorithm run and then reassigned to the other fixed value for a second subsequent run. This assignment ensures that we can infer the behavior of $\mathcal{T}(S_{n})$ based on the algorithm’s output. To do this, we need to construct and evaluate a two-part unitary function $U_{f}$ across two distinct runs of the algorithm.
@@ 5.1.1 Behavior for $S_{n}$
For $s_{i}\in S_{n}$, the oracle function evaluates $\mathcal{T}(s_{i})$, flipping the phase of |$\ket{s_{i}}$ when $\mathcal{T}(s_{i})=1$ and leaving it unchanged when $\mathcal{T}(s_{i})=0$, such that
$\text{If }\mathcal{T}(s_{i})=1:\quad U_{f}\ket{s_{i}}\ket{-}=-\ket{s_{i}}\ket{-},$ (27a)
$|.\displaystyle\text{If }\mathcal{T}(s_{i})=0:\quad U_{f}\ket{s_{i}}\ket{-}=\ket{s_{i}}\ket{-}.$ (27b)
@@ 5.1.2 Behavior for $S_{n}^{c}$
For $s_{i}\in S_{n}^{c}$, the oracle does not evaluate $\mathcal{T}(s_{i})$ directly. Instead:
@@
    For the first run of the algorithm, all elements $s_{i}\in S_{n}^{c}$ are assigned a value of 0, i.e., $\mathcal{T}(S_{n}^{c})=0$ always.
@@
    In a separate second run, all elements $s_{i}\in S_{n}^{c}$ are assigned a value of 1, i.e., $\mathcal{T}(S_{n}^{c})=1$ always..
This ensures that $S_{n}^{c}$ does not contribute to the dynamic information encoded in $\mathcal{T}(S_{n})$. But it satisfies the technical requirement that any input function $f(\cdot)$ into Deutsch-Jozsa operates over the entire domain $\mathbb{S}_{n}$ by filling in the values of $\mathbb{S}_{n}$ not encountered in $S_{n}$ that encodes the dynamics.
@@
$\text{If assigned }\mathcal{T}(s_{i})=1:\quad U_{f}\ket{s_{i}}\ket{-}=-\ket{s_{i}}\ket{-},$ (28a)
$|.\displaystyle\text{If assigned }\mathcal{T}(s_{i})=0:\quad U_{f}\ket{s_{i}}\ket{-}=\ket{s_{i}}\ket{-}.$ (28b)
@@ 5.2 Constructing $U_{f}$ to Evaluate $f(S_{n})$
    Evaluation of $S_{n}$: Apply $\mathcal{T}(s_{i})$ to elements in $S_{n}$ according to equation 27.
@@
The second step evaluates $\mathcal{T}(S_{n})$ by comparing each $\Sigma_{r}$ (encoded in $s_{i}$) with the threshold $\Sigma_{T}$. We implement this as a binary subtraction $\Sigma_{r}-\Sigma_{T}$, with the ancillary qubit encoding the result
@@
    |$\ket{1}$ if $s_{i}\in S_{n}=\Sigma_{r}\geq\Sigma_{T}$, corresponding to $\mathcal{T}(s_{i})=1$.
@@
    |$\ket{0}$ if $s_{i}\in S_{n}=\Sigma_{r}<\Sigma_{T}$, corresponding to $\mathcal{T}(s_{i})=0$.
@@ 5.2.1 Example of a Binary Subtraction for Evaluating $f(S_{n})$
Let’s work through a specific example of a binary subtraction operation. Consider a vertex $v_{j}\in G(V,E)$ with a corresponding entry $\nu_{i}\in\mathbb{V}_{n}$ that encodes the running summation value $\Sigma_{r}$ at the observation time $T_{o}$ for the specific vertex $v_{j}$. This value would be marked as an element $s_{i}\in S_{n}$. Assume a 7-bit representation of $\Sigma_{r}$ and the threshold $\Sigma_{T}$ to two-decimal precision, as discussed in Section 4.1.
@@
where |$\ket{\nu}$ encodes $\Sigma_{r}=0.85$ in binary, while |$\ket{\tau}$ encodes $\Sigma_{T}=0.50$. The ancillary qubit, initially in state |$\ket{0}$, will store the result of comparing $\Sigma_{r}\geq\Sigma_{T}$.
@@
This binary subtraction comparison is efficient because it focuses only on detecting the sign of the result (i.e., positive or negative) rather than computing the full numerical value of $\Sigma_{r}-\Sigma_{T}$. We discuss the thresholding sub-circuit that we used in our numerical simulations and experiments to implement this scheme in section 7.3.
5.3 Measurement Interpretations of $U_{f}$
By running the Deutsch–Jozsa algorithm twice, assigning all elements of $S_{n}^{c}$ first to $f(s_{i})=0$ and then to $f(s_{i})=1$, measurement outcomes provide an interpretation that answers our question about the network dynamics (Box 1.):
@@ 5.3 Measurement Interpretations of $U_{f}$
    Quiescence: If $S_{n}^{c}=0$ and $U_{f}$ evaluates to a constant state |$\ket{1}^{\otimes n}$, the network is quiescent, because this can only occur when $\mathcal{T}(S_{n})=0$ for all $s_{i}\in S_{n}$.
@@
    Epileptic State: If $S_{n}^{c}=1$ and $U_{f}$ evaluates to a constant state |$\ket{0}^{\otimes n}$, the network is epileptic, because this can only occur $\mathcal{T}(S_{n})=1$ for all $s_{i}\in S_{n}$.
@@
    Dynamic Activity: If $U_{f}$ is not constant for both cases, we interpret this result as the network being able to potentially maintain inherent dynamic activity due to the measurement outcome reflecting what must be a mix of active and inactive vertices in $S_{n}$.
@@ 6 Comparative Algorithmic Complexity
In this section, we analyze the algorithmic complexity of solving the problem using our workflow leveraging the Grover and Deutsch–Jozsa algorithms versus a purely classical approach. We will show that using these quantum algorithms theoretically conveys a significant computational advantage and speedup and demonstrate how such an advantage scales, i.e., becomes greater, as the size of the network being analyzed increases.
@@ 6.1 Classical Complexity
    Evaluate $\mathcal{V}$ for each $v_{j}\in V$ in order to compare $\Sigma_{r}$ and $\Sigma_{T}$ (equation 16). This requires iterating over all vertices.
@@
    Combine the results of for all $v_{j}\in V$ to determine whether the network is quiescent, epileptic, or dynamically active.
@@ 6.2 Quantum Complexity: Grover’s Algorithm
Grover’s algorithm provides a quadratic speedup for searching an unstructured set. To determine $S_{n}$, we use Grover’s algorithm to find all $s_{i}\in\mathbb{S}_{n}$ that belong to $S_{n}$. Let $\mu$ denote the number of occurrences of a particular $s_{i}$ corresponding to a a value $\Sigma_{r}\in\mathbb{V}_{n}$, and $|\mathbb{S}_{n}|=2^{n}$. The number of iterations required for Grover’s algorithm to amplify the amplitude of each solution is approximately
@@ 6.3 Quantum Complexity: Deutsch–Jozsa Algorithm
The Deutsch–Jozsa algorithm evaluates whether $\mathcal{T}(S_{n})$ is constant or not across all elements of $\mathbb{S}_{n}$. Unlike a classical approach that would require checking each $s_{i}\in\mathbb{S}_{n}$, the Deutsch–Jozsa algorithm achieves this in a constant number of evaluations:
@@ 6.4 Comparative Analysis
As such, the task of evaluating $\mathcal{T}(S_{n})$ for all vertices in $V$ can be summarized by
@@
For large $N$ and $n$, a quantum algorithmic approach provides a significant increase in computational efficiency. Grover’s algorithm provides a quadratic speedup in $2^{n}$, while Deutsch–Jozsa evaluates $\mathcal{T}(S_{n})$ in constant time.
@@
Subsequently evaluating $\mathcal{T}(S_{n})$ with Deutsch–Jozsa adds a constant overhead of $O_{\text{Deutsch-Jozsa}}=O(1).$
@@
We can assess the comparative theoretical speedup advantage of the quantum approach versus the classical approach by computing the ratios of their respective complexities. The classical approach needs $O(7\times 10^{6})$, while the quantum approach requires $O(1131)$. This then translates to a speedup factor of approximately
@@
By ‘faster’, we specifically mean that a quantum approach, the way we describe it here, requires significantly fewer computational steps to arrive at the result. This reduction translates into less computational time, assuming similar hardware clock speeds and negligible algorithmic overhead. This also assumes a physical quantum computer with sufficiently stable logical qubits.